# Solving rescheduling problems in heterogeneous urban railway networks using hybrid quantum-classical approach

### v5

@@ 4.2.2 Constraints
These ensure that trains cannot meet in interlocking areas:
@@ 4.2.3 Scaling analysis
Let us discuss the number of variables and constraints in our model. The number of time variables #$\#t$, the product #$\#\mathcal{J}\#\mathcal{S}$ of the number of trains and the number of stations serves as a trivial but rough upper bound. (See also Section 3.1 of the work of  [17]). To obtain a more accurate estimate of the number of these variables we observe that in our model no train visits all the stations. To provide an estimate, we introduce $\alpha\in(0,1)$ as the fraction of the stations visited by an average train. Thus we have
#$\#(t)\approx\alpha\#\mathcal{J}\#\mathcal{S}\leq\#\mathcal{J}\#\mathcal{S}.$ (20)
as a tighter bound. In what follows we will mostly $\alpha=1$, the rough bound, however, we will keep the parameter $\alpha$ in the formulae to reflect its role in the scaling. Let us assume, that half the of trains (#$\frac{\#\mathcal{J}}{2}$) are going in one direction and the other half are going in the other. Then the number of trains each one meets going in the same direction is $\alpha\frac{\#\mathcal{S}\#\mathcal{J}}{2}$. This yields #$\#\mathcal{J}\cdot\frac{\alpha\#\mathcal{S}\#\mathcal{J}}{2}$ $y^{(\text{out})}$ variables. However, because of the symmetry property in Equation (19), only half of these variables are independent, thus we have $\alpha\#\mathcal{S}\frac{\#\mathcal{J}^{2}}{4}$ $y^{(\text{out})}$ of these variables. As for the trains going in opposite directions, we assume that each train meets at most #$\frac{\#\mathcal{J}}{2}$ trains on each station in the single-track parts of the network. Hence, the number of $z$ variables is at most #$\#\mathcal{J}\cdot\frac{1}{2}\cdot\frac{\alpha\#\mathcal{J}(\#\mathcal{S}-1)}{2}$ (here, #$\#\mathcal{S}-1$ comes from the observation that the $z$-variables are tied to pairs of stations). Variables $y^{(\text{in})}$ occur mainly in Eqs. (4.2.2)-(4.2.2), relevant for a few stations only: there is typically one such station in the network, yielding at most #$\frac{\#\mathcal{J}^{2}}{4}$ $y^{(\text{in})}$ variables. In summary, the number of precedence variables can be estimated as
#$\begin{split}&\#y^{(\text{out})}\leq\alpha\frac{\#\mathcal{J}^{2}}{4}\#\mathcal{S}\leq\frac{\#\mathcal{J}^{2}}{4}\#\mathcal{S}\\ &\#y^{(\text{in})}\leq\frac{\#\mathcal{J}^{2}}{4}\\ &\#z\leq\alpha\frac{\#\mathcal{J}^{2}}{4}(\#\mathcal{S}-1)\leq\frac{\#\mathcal{J}^{2}}{4}(\#\mathcal{S}-1)\\ &\#y+\#z=\#y^{(\text{out})}+\#y^{(\text{in})}+\#z<\frac{\#\mathcal{J}^{2}}{2}\#\mathcal{S}\end{split}$ (21)
@@
Note that these bounds are rather pessimistic: setting the parameter $d_{\max}$ as an upper bound on the secondary delays implies a reduction in the number of trains that can potentially conflict. However, the actual number of them can depend on the train density and structure of the timetable in various ways, hence, it is not possible to provide a tighter bound that would be valid in general. Finally, let us remark as the $z$-variables are mainly tied to single-track traffic, the model is more complex for networks dominated by single-track segments.
Based on Eqs. (20) and Equation (21), the scaling of the upper bound on the the total number of decision variables can be estimated by
#$\#vars=\#t+\#y+\#z\approx\text{const}\#\mathcal{J}^{2}\#\mathcal{S}+\text{const'}\#\mathcal{J}\#\mathcal{S},$ (22)
@@
Now let us estimate the number of constraints. Considering the headway constraints in Equation (4), there is one such constraint for each variable. However, because of the symmetry property in Equation (19), only half of the variables are independent, hence, the number of these constraints is twice the number of independent variables.
Along the same arguments, in case of the constraints in Equation (11), i.e. station track occupancies, there are typically $2$ of these for each independent $y^{(\text{out})}$ variable, and for the single-line track occupancy constraints in Equation (9) there are $2$ constraints per each $z$ variable. It is more complicated to estimate the number of interlocking area occupation constraints in Eqs. (4.2.2))-(4.2.2)) as they involve $z$, $y^{(\text{out})}$ and $y^{(\text{in})}$ variables. A good assumption is to consider $2$ constraints for each $y^{(\text{out})}$ variable, $2$ constraints for each $y^{(\text{in})}$ variable, and $2$ constraints for each $z$ variable. Further, we expect one minimal dwell time constraint per train and station (described by a $\alpha\#\mathcal{J}\#\mathcal{S}$ term). Concerning train order constraints in Eqs (17) (18), we also expect one constraint per train and station (also described by a $\alpha\#\mathcal{J}\#\mathcal{S}$ term). We exclude the rolling stock circulation constraints from this consideration as their contribution to the number of constraints is very limited. In summary, the total number of constraints obeys the following upper bound:
#$\#constr.<6\#y^{(\text{out})}+2\#y^{(\text{in})}+4\#z+2\alpha\#\mathcal{J}\#\mathcal{S}.$ (23)
Substituting the estimated number of variables in Equation (21) we have
#$\#constr.<\alpha\frac{3}{2}\#\mathcal{J}^{2}\#\mathcal{S}+\frac{1}{2}\#\mathcal{J}^{2}+\alpha\#\mathcal{J}^{2}(\#\mathcal{S}-1)+2\alpha\#\mathcal{J}\#\mathcal{S}$ (24)
Setting $\alpha=1$ as the worst case, we get
#$\#constr.<\frac{5}{2}\#\mathcal{J}^{2}\#\mathcal{S}-\frac{1}{2}\#\mathcal{J}^{2}+2\#\mathcal{J}\#\mathcal{S}.$ (25)
@@ 7 Conclusions
We have demonstrated that quantum annealers can be readily applicable in train rescheduling optimization in urban railway networks. In particular, we have encoded rescheduling problems in a linear integer program and solved these using a state-of-the-art solver (CPLEX), as well as the QA based CQM hybrid (quantum-classical) solver. The results suggest that quantum computing and QA in particular, although an early-stage technology, are ready to tackle challenging railway rescheduling problems. While the CQM hybrid solver does not outperform classical solvers in general, we have found specific cases in which they were competitive. This supports the expectation that future quantum devices will be efficient in solving large-scale problems, and in our case, real-time railway dispatching problems on the national scale; even in the range that is beyond the scope of current exact models and heuristics.
While the quantum-based solvers possibly return suboptimal solutions, these are still feasible and often better than the solutions obtained manually or based on smaller-scale models. In each computation, the CQM hybrid solver always has actually used the QPU. Hence, such a little help from the QPU may have boosted the classical heuristics in the hybrid solver, even if quantum annealer devices are subjects of many ongoing debates.
Several future research directions can be determined. The first concerns the ongoing worldwide works on the development of fault-tolerant quantum computing devices [13] with proper error correction/mitigation, this will boost results of approaches supported by quantum computing (and QA in particular). The second concerns the creation of a custom open-source hybrid solver that is dedicated to railway problems and will use the noise of the quantum device to our advantage e.g. by modeling some stochastic behavior on railways. The third concerns the evolution and application of quantum-inspired techniques such as tensor networks technique, and simulated bifurcation [20]. With all these future developments, QA-based approaches tend to become more powerful and ready to address for the first time large-scale real-time challenges in railway traffic management, which are currently unsolvable with traditional techniques.

### v6

@@ 4.2.2 Constraints
These ensure that trains that use the same switch cannot meet in interlocking areas:
@@ 4.2.3 Scaling analysis
Let us discuss the number of variables and constraints in our model. To assess the problem’s size we will use here a worst-case approximation, namely, we assume temporarily that all trains visit all stations and any pair of trains can meet each other; these will be assumed throughout the entire consideration on scaling. Under such assumptions the number of time variables #$\#t$ can be upper bounded by the product #$\#\mathcal{J}\#\mathcal{S}$ of the number of trains and the number of stations:
#$\#(t)\leq\#\mathcal{J}\#\mathcal{S}.$ (20)
Each train meets at most (#$(\#\mathcal{J}-1)\#\mathcal{S}$ trains going in same direction, which results in #$\#\mathcal{J}(\#\mathcal{J}-1)\#\mathcal{S}$ $y^{(\text{out})}$ variables in the model. However, because of the symmetry property in Equation (19), only half of these variables are independent, thus we have at most #$\frac{\#\mathcal{J}(\#\mathcal{J}-1)\cdot\#\mathcal{S}}{2}$ of these variables. As for the trains going in opposite directions, each train meets #$\#\mathcal{J}-1$ trains at each station. (Although we have a quadruple of indices $(j,j^{\prime},s,s^{\prime})$ in the definition of $z$ variable, as routes of trains are fixed within each model instance, the elements of this quadruple are not independent. In particular, the index $s^{\prime}$ is uniquely defined by the other indices). Hence, the number of $z$ variables is at most #$\frac{\#\mathcal{J}(\#\mathcal{J}-1)\cdot\#\mathcal{S}}{2}$ (having taken into account the symmetry property in Equation (19) again). Variables $y^{(\text{in})}$ occur mainly in Equations (4.2.2)-(4.2.2), and are relevant relevant for some stations only. Nevertheless, in the worst case there can be one such variable assigned for each pair of trains and each station, yielding #$\frac{\#\mathcal{J}(\#\mathcal{J}-1)\cdot\#\mathcal{S}}{2}$ $y^{(\text{in})}$ variables after the reduction by symmetry properties. In summary, the upper bound for the number of precedence variables reads
#$\begin{split}\#y^{(\text{out})}&\leq\frac{\#\mathcal{J}(\#\mathcal{J}-1)\cdot\#\mathcal{S}}{2}\\ \#y^{(\text{in})}&\leq\frac{\#\mathcal{J}(\#\mathcal{J}-1)\cdot\#\mathcal{S}}{2}\\ \#z&\leq\frac{\#\mathcal{J}(\#\mathcal{J}-1)\cdot\#\mathcal{S}}{2}\\ \#y+\#z&=\#y^{(\text{out})}+\#y^{(\text{in})}+\#z\leq\frac{3}{2}\#\mathcal{J}(\#\mathcal{J}-1)\#\mathcal{S}\end{split}$ (21)
@@
Note that these bounds are rather pessimistic: setting the parameter $d_{\max}$ as an upper bound on the secondary delays implies a reduction in the number of trains that can potentially conflict. However, the actual number of them can depend on the train density and structure of the timetable in various ways, hence, it is not possible to provide a tighter bound that would be valid in general. Finally, let us remark that as the $z$-variables are mainly tied to single-track traffic, the model is more complex for networks dominated by single-track segments. Based on Equations (20) and (21), the scaling of the upper bound on the total number of decision variables can be estimated by
#$\#vars=\#t+\#y+\#z\approx O(\#\mathcal{J}^{2}\#\mathcal{S}),$ (22)
@@
Now let us estimate the number of constraints. Considering the headway constraints in Equation (4), there are two such constraints for each independent $y^{(\text{out})}$ variable. (Because of the symmetry property in Equation (19), only half of the $y^{(\text{out})}$ variables are independent).
Along the same arguments, in case of the constraints in Equation (11), i.e. station track occupancies, there are typically $2$ of these for each independent $y^{(\text{out})}$ variable, and for the single-line track occupancy constraints in Equation (9) there are $2$ constraints per each independent $z$ variable. It is more complicated to estimate the number of interlocking area occupation constraints in Equations (4.2.2))-(4.2.2)) as they involve $z$, $y^{(\text{out})}$ and $y^{(\text{in})}$ variables. A good assumption is to consider $2$ constraints for each independent $y^{(\text{out})}$ variable, $2$ constraints for each independent $y^{(\text{in})}$ variable, and $2$ constraints for each independent $z$ variable. Further, we expect one minimal dwell time constraint per train and station, resulting in #$\#\mathcal{J}\#\mathcal{S}$ constraints in the worst case. Concerning train order constraints in Equations (17) and (18), there can be one constraint per train and station, which yields additional #$\#\mathcal{J}\#\mathcal{S}$ constraints in the worst case. As we do not consider splitting and joining trains on the route, at most one rolling stock circulation constraint arises per train, yielding at most #$\#\mathcal{J}$ such constraints. In summary, the total number of constraints obeys the following (worst-case) upper bound:
#$\#constr.\leq 6\#y^{(\text{out})}+2\#y^{(\text{in})}+4\#z+2\#\mathcal{J}\#\mathcal{S}+\#\mathcal{J}.$ (23)
Substituting the estimated number of variables in Equation (21) we have:
#$\#constr.\leq 6\#\mathcal{J}^{2}\#\mathcal{S}-4\#\mathcal{J}\#\mathcal{S}+\#\mathcal{J}.$ (24)
@@ 7 Conclusions
We have demonstrated that quantum annealers can be readily applicable in train rescheduling optimization in urban railway networks. In particular, we have encoded rescheduling problems in a linear integer program and solved these using a state-of-the-art solver (CPLEX), as well as the QA based CQM hybrid (quantum-classical) solver. The results suggest that quantum computing and QA in particular, although an early-stage technology, are ready to tackle challenging railway rescheduling problems. While the CQM hybrid solver does not outperform classical solvers in general, we have found specific cases in which they were competitive. These results support the expectation that future quantum devices can become efficient in solving large-scale problems, like, in our case, real-time railway dispatching problems on the national scale, even in the range that is beyond the scope of current exact models and heuristics.
While the quantum-based solvers possibly return suboptimal solutions, these are still feasible and often better than the solutions obtained manually or based on smaller-scale models. It is important to note that the CQM hybrid solver has always used the QPU in each computation. The actual assistance from the QPU may have boosted the classical heuristics in the hybrid solver.
A possible way of improving on the presented technique would be to find a way of determining the smallest suitable $d_{\text{max}}$ based on the input data instead of an ad-hoc choice, by using some fast enough methods. Apart from that, several future research directions can be determined. The first concerns the ongoing worldwide efforts on the development of fault-tolerant quantum computing devices [13] with proper error correction/mitigation. Those future quantum devices are expected to boost the efficiency of approaches supported by quantum computing (and QA in particular). The second concerns the creation of a custom open-source hybrid solver that is dedicated to railway problems and will use the noise of the quantum device to our advantage e.g. by modeling some stochastic behavior on railways. The third concerns the evolution and application of quantum-inspired techniques such as tensor network techniques or simulated bifurcation [20]. With all these future developments, QA-based approaches tend to become more powerful and ready to address for the first time large-scale real-time challenges in railway traffic management, which are currently unsolvable with traditional techniques.